One quarter of the sweets manufactured for packets ofv"Lucky Five" were pink. Each packet contains five sweets. Out of 1024 packets how many would you expect to contain at least two pink sweets?
1
Expert's answer
2019-04-09T09:57:46-0400

"P(X\u22652)=1-P(X<2)=1-P(X=0)-P(X=1)=\n1-C_5^0 0.25^0 0.75^5-C_5^1 0.25^1 0.75^4=0.3672.\n\nN=1024*0.3672=376."